Fermi National Laboratory


Accelerator Update

Friday February 20
The day shift began with Operations monitoring store 3245 and the Recycler stash, and with the Antiproton Source (Pbar) stacking.

At 8:15 AM, Tevatron (TeV) experts began a Synclight (beam position) study.

Operations terminated store 3245 at 9:11 AM.

Operations put the TeV into shot setup at 10:07 AM. The Recycler will use its stash to establish this store.

Operations began sending beam to MiniBooNE at 10:47 AM.

At 12:38 PM, Operations established store 3247, from the Recycler, with an initial luminosity of 18.85E30. Pbar resumed stacking and TeV experts began separator studies.

A Main Injector (MI) expert began a slipstacking study at 1:39 PM.

A Switchyard (SY) expert began a study period at 1:53 PM. The beam was being sent to the SY dump.

At 3:55 PM, a Recycler expert took protons from MI.

Both TeV and SY experts had completed their studies by 4:36 PM.

MI studies finished at 5:05 PM.

At 9:04 PM, Operations terminated store 3247.

At 23:59 PM, Operations established store 3249 with an initial luminosity of 51.5E30.

Saturday February 21
The midnight shift began with Operations monitoring store 3249, with Pbar stacking, and with MiniBooNE taking beam.

At 1:17 AM, the Pbar stacking rate dropped by 50%. Operators investigated and found a problem with the Debuncher septa (D:ESEPV). They called in an expert.

The septa expert reported at 4 AM that he had found a bad cable connection. Unfortunately, the cable was too short to re-terminate, so he had to make a replacement cable. The expert had this new cable installed by 5:15 AM. Pbar resumed stacking.

A Recycler expert began stashing shots of protons from MI at 9:27 AM.

Booster had been suffering from beam problems that operators fixed by tuning at 3:57 PM.

Sunday February 22
The midnight shift began with Operations monitoring store 3249, with Recycler taking shots of protons, with Pbar stacking, and with MiniBooNE taking beam.

At 4:05 AM, Operations terminated store 3249.

Operations put the TeV into shot setup and established store 3252 at 8:10 AM. Its initial luminosity was 57.03E30.

A SY expert began a six-hour study at 3:51 PM.

A MI expert began a study at 5:43 PM.

Monday February 23
The midnight shift began with Operations monitoring store 3252, with MiniBooNE taking beam, and with Pbar stacking.

At 12:48 AM, while conducting a tour of TeV areas, operators discovered a blown fuse for a heater. They called the Duty Electrician who locked it off for later repairs.

Today's Plans
The plans for today are to continue stacking and then later this afternoon establish a new store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/23/2004   email Fermilab